OpenStudy (anonymous):

Find the difference. 4d2e - d2e. Select one of the options below as your answer: A. 3 B. 3de C. 3d2e D. 4d2e

OpenStudy (anonymous):

the 2 is squared on both numbers

OpenStudy (anonymous):

4d2e - d2e we have same things caz they have equal powers(1) and same worlds(d&e) so do this : 4d2e - d2e=(4-1)d^2e=3d^2e Got it ?:)

OpenStudy (anonymous):

C. 3d2e :) Mathway.com helps a lot too lol